The perimeter of a field is 140 feet. The width is one-sixth of the length. Find the width of the field.
Answer by Maths68(1474)

Perimeter of a field = P = 140 feet
Length = L
Width = W=1/6(L)=L/6
P = 2L + 2W
Put values in above equation
140=2L+2(L/6)
140=2L+2L/6
140=2L+L/3
Multiply by 3 both sides
420=6L+L
420=7L
420/7=L
L=60
Given
W=L/6
W=60/6
W=6feet